### v2

# Planckian bound on quantum dynamical entropy

Abstract
We introduce a simple definition of dynamical entropy for quantum systems under continuous monitoring, inspired by Connes, Narnhofer and Thirring. It quantifies the amount of information gained about the initial condition. A nonzero entropy rate can be obtained by monitoring the thermal fluctuation of an extensive observable in a generic many-body system (away from classical or large $N$ limit). We explicitly compute the entropy rate in the thermodynamic and long-time limit, in terms of the two-point correlation functions. We conjecture a universal Planckian bound for the entropy rate. Related results on the purification rate are also obtained.
Introduction. The dynamical entropy of Kolmogorov and Sinai [1, 2, 3] is an important concept in classical mechanics and thermodynamics. It can be defined as the amount of information that an observer may gain about the initial condition by continuously monitoring the system. The dynamical entropy is a fundamental diagnosis of classical chaos [4]. The butterfly effect amplifies arbitrarily small details of the initial condition, and makes them observable at late time. This leads to a nonzero dynamical entropy rate, which is quantitatively tied to the Lyapunov exponents by Pesin’s relation [5]. Relations have been established with thermodynamic entropy production [6] and transport [6, 7].
Extending the notion of dynamical entropy to quantum systems has been an outstanding question since Kolmogorov and Sinai’s work [8]. A few different proposals have been put forward by several groups [9, 10, 11, 12, 13, 14, 15, 16, 17]. Unfortunately, these works have received limited attention from physicists. Concrete calculations are often difficult and only performed for simple systems/dynamics such as free gas or spatial translation [10, 18, 14, 19]. More recently, the question of describing chaos and complexity in quantum many-body systems has been intensively pursued from various viewpoints [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. Yet it remains unclear whether some notion of quantum dynamical entropy can be valuable in this context (see [36] for a recent survey.)
In this Letter, inspired by the work of Connes, Narnhofer and Thirring (CNT) [10], we consider a simple definition of quantum dynamical entropy via continuous monitoring  [19, 37, 38]. We focus on quantum many-body systems with local interaction, in which one monitors a “mesoscopic observable”, which is an appropriately rescaled sum of local operators, see (10)-(11) below. The main result is an exact formula (13) that computes the asymptotic dynamical entropy rate in the thermodynamic limit, when the initial state is a thermal equilibrium. The rate is given in terms of the thermal correlation function. The formula applies generally, with the exception of thermal critical points. When it applies, the entropy rate obeys a Planckian bound: in units with $\hbar=k_{B}=1$,
$\lim_{t\to\infty}\lim_{V\to\infty}S_{\text{CNT}}/t\leq cT\,$, (1)
where $T$ is the temperature and $c$ is a universal constant, see (14) below. This bound is reminiscent of that on the growth rate out-of-time order correlation functions (OTOCs) [39] (see [40, 41, 42, 43, 44, 45] for other “Planckian bounds”). OTOCs generalize the butterfly effect in quantum systems, and have an extended exponential growth in semiclassical or large $N$ systems (among which the Sachdev-Ye-Kitaev models saturate the Planckian bound [46, 47, 48, 49, 50]). The dynamical entropy studied here has a well-defined linear growth in generic quantum many-body systems, away from semiclassical or large $N$ limit. This feature and the Planckian bound suggest the dynamical entropy as an interesting probe of low-temperature many-body quantum chaos.
We obtain similar results for the amount by which the monitoring purifies a thermal initial state. This quantity is akin to the dynamical entropy, which also uses the idea of purification in a crucial way. Recently, purification has been studied as a probe of measurement-induced phases of information  [51, 52, 53, 54, 55]. Under sufficiently weak monitoring, the long time regime of complete purification is known to admit a universal description [56, 57]. Our results are complementary, and capture model dependence at short-time in a general way.
Setup. The dynamical entropy is defined using a setup of purification by continuous monitoring, which we recall briefly here, see Fig. 1 for illustration.
Figure 1: The purification setup. Two identical systems $A\bar{A}$ are initialized in an entangled state $\sqrt{\rho}$. AA is evolved by a unitary $U$ and repeatedly measured, yielding outcomes $m_{\delta t},\dots,m_{t}$. The latter disentangles (purifies) $\bar{A}$, revealing information about the initial condition.
Consider a quantum system $A$, whose initial state is mixed, and given by the density matrix $\rho$. We shall purify $\rho$ as the reduced density matrix of the pure state $\sqrt{\rho}$ on $A\bar{A}$, where $\bar{A}$ is an identical copy of $A$. (The reduced density matrix of $\bar{A}$ is also $\rho$.) We shall by default be concerned with the case where $A$ is a many-body quantum system with short-range Hamiltonian $H$, and $\rho=\frac{1}{Z}e^{-\beta H}$ is the thermal equilibrium at temperature $1/\beta$. In that case $\sqrt{\rho}$ is known as the thermofield double (see [58, 59] for realization proposals on near-term quantum platforms).
The system $A$ undergoes discrete time evolution governed by a unitary operator $U$, whereas $\bar{A}$ is left intact. By default, $U=e^{-iH\delta t}$, and we will eventually take the continuous-time limit $\delta t\to 0$. After every time step, we perform a same measurement on $A$, described by the Kraus operators $\{K_{m}\}_{m}$ [60], indexed by the possible outcomes $m$, satisfying $\sum_{m}K_{m}^{\dagger}K_{m}=\mathbf{1}$ (unitarity). Doing this until $t=n\delta t$, we obtain a random sequence of outcomes $\vec{m}=(m_{\delta t},m_{2\delta t},\dots,m_{t})$. In terms of
$\tilde{\rho}_{\vec{m}}:=\sqrt{\rho}\prod_{s=t}^{\delta t}K_{m_{s}}^{\dagger}(s)\prod_{s=\delta t}^{t}K_{m_{s}}(s)\sqrt{\rho}\,$,\, (2)
where $O(k\delta t):=U^{-k}OU^{k}$, the outcome probability and the post-measure state of $\bar{A}$ are respectively
$p_{\vec{m}}=\mathrm{Tr}[\tilde{\rho}_{\vec{m}}]\,,\,\rho_{\vec{m}}=\tilde{\rho}_{\vec{m}}/p_{\vec{m}}\,.$ (3)
The conditional entropy is defined as the measurement-induced decrease of the von Neumann entropy of $\bar{A}$, averaged over the outcomes (weighed by Born’s rule):
$J:=\sum_{\vec{m}}p_{\vec{m}}\mathrm{Tr}[\rho_{\vec{m}}\ln\rho_{\vec{m}}-\rho\ln\rho]\,.$ (4)
$J$ can be viewed as the mutual information between $\bar{A}$ and the (dephased) measurement ancillas, and quantifies the amount of initial condition revealed by the observer [61, 62]. We shall be interested in the purification rate $J/t$ in the long time limit.
The definition of our CNT inspired dynamical entropy is slightly more involved. For each $s=\delta t,\dots,t$, we introduce the marginal analogues of (2) and (3)
$\tilde{\rho}^{s}_{m}:=\sum_{\vec{m}:m_{s}=m}\tilde{\rho}_{\vec{m}}\,,\,p^{s}_{m}:=\mathrm{Tr}[\tilde{\rho}^{s}_{m}]\,,\,\rho^{s}_{m_{s}}:=\frac{\tilde{\rho}^{s}_{m}}{p^{s}_{m}}\,$, (5)
and the marginal analogue of (4):
$J_{s}:=\sum_{m_{s}}p_{m_{s}}^{s}\mathrm{Tr}[\rho^{s}_{m_{s}}\ln\rho_{m_{s}}^{s}-\rho\ln\rho]\,.$ (6)
We are ready to define the quantum dynamical entropy:
$S_{\text{CNT}}:=$ $\sum_{\vec{m}}-p_{\vec{m}}\ln p_{\vec{m}}+\sum_{s=\delta t}^{t}\left[\sum_{m_{s}}p^{s}_{\vec{m}}\ln p^{s}_{\vec{m}}+J_{s}\right]$
$=$ $S_{\text{cl}}(\{p_{\vec{m}}\})-\sum_{s=\delta t}^{t}\left[S_{\text{cl}}(\{p^{s}_{m}\})-J_{s}\right]\,.$ (7)
Namely, it is defined as the classical Shannon entropy of the joint outcome distribution $S_{\text{cl}}(\{p_{\vec{m}}\})$, minus a sum of “entropy defects”, $S_{\text{cl}}(\{p^{s}_{m_{s}}\})-J_{s}$, which are the difference between the outcome Shannon entropy and the information gain of a single measurement.
The entropy defect was introduced by CNT [10] as the key ingredient to extend the Kolmolgorov-Sinai entropy to quantum systems. It accounts for a key feature of quantum physics: measurement inevitably disturbs the system, and thus can prevent later observations from gaining information about the initial state [19, 16]. Such disturbance results in a lower quantum dynamical entropy (Planckian bound on quantum dynamical entropy). This point will be important as we discuss different types of monitoring below.
The dynamical entropy (Planckian bound on quantum dynamical entropy) depends on the duration ($t$), dynamics ($U$), the initial state ($\rho$), and the measurement scheme $\{K_{m}\}_{m}$. We will call $S_{\text{CNT}}/t$ the entropy rate, and be interested in its limit as we first take the thermodynamic limit and then $t\to\infty$.
Our definition (Planckian bound on quantum dynamical entropy) has the same structure as CNT’s proposal. Technically, in that work, the measurements are replaced by a more general notion of “abelian models”, over which a supremum is then taken. Here, we do not take a supremum over the measurement scheme. Indeed, in a many-body system, its choice is a nontrivial matter, that we discuss next.
Choice of observable. To start, we shall restrict to monitoring a single real-valued observable. In particular, we do not measure an extensive number of sites at each time step. Doing so substantially alters the intrinsic dynamics, which we would like to probe using dynamical entropy. This restriction is also necessary for a Planckian bound, as the rate can be trivially increased by monitoring multiple weakly correlated regions.
Within this restriction, a natural possibility is a microscopic measurement, which acts on $O(1)$ local degrees of freedom. Note that we assume those to be non-semiclassical, and have a small Hilbert space. We argue that the asymptotic CNT entropy rate cannot be positive for a generic dynamics, due to the significant disturbance caused by microscopic measurements.
In fact, we shall argue that $J_{t}\to 0$ as $t\to\infty$; from the definition (Planckian bound on quantum dynamical entropy) and the subaddictivity of the Shannon entropy, we have $S_{\text{CNT}}/t\leq\sum_{s\leq t}J_{s}/t$, so $J_{t}\to 0$ would forbid a positive entropy rate. For that, observe from (5), (2), and unitarity, that
$\tilde{\rho}^{t}_{m}=\sqrt{\rho}O_{m}\sqrt{\rho}\,,\,O_{m}=\mathcal{L}(\mathcal{D}\mathcal{L})^{n-1}[K_{m}^{\dagger}K_{m}]\,.$ (8)
Here $n=t/\delta t$ and
$\mathcal{L}[O]:=U^{\dagger}OU\,,\,\mathcal{D}[O]:=\sum_{m}K_{m}^{\dagger}OK_{m}\,$,\, (9)
are the time-evolution and dephasing channel (induced by a measurement with discarded result). For concreteness, consider strongly measuring a spin in a lattice along the $z$ direction, then $K_{m=\pm}=(1\pm Z)/\sqrt{2}$ ($Z$ is the Pauli acting on the spin), and $\mathcal{D}[(aX+bY)O^{\prime}]=0$ for any $a,b$, where $X$ or $Y$ acts on the monitored spin and $O^{\prime}$ is any operator acting elsewhere. Such terms are repeatedly generated by a generic enough $\mathcal{L}$ and annihilated by $\mathcal{D}$, until the operator $(\mathcal{D}\mathcal{L})^{n}[K_{m}^{\dagger}K_{m}]\to c\mathbf{1}$ as $n\to\infty$. Thus, $\tilde{\rho}^{t}_{m}$ tends to be $\propto\rho$ and $J_{t}\to 0$.
Figure 2: Exact brute-force numerical calculation of the quantity $J_{s}$ (6) and the purification rate $J/t$ (4) with microscopic (a,c) vs mesoscopic (b,d) monitoring ($\delta t=1$), on mixed-field Ising chains of size $L=6,8,10,12$s. $H=\sum_{j=1}^{L}X_{j}+hZ_{j}+JZ_{j}Z_{j+1}$ ($L+1\equiv 1$), $h=1.245,J=0.945$ at $\beta=0$. The Kraus operators are $K_{m=\pm}=\sqrt{(\mathbf{1}\pm Q)/2}$ with $Q=\gamma Z_{0}Z_{2}$ in (a,c) and $Q=\mathrm{erf}\left(\gamma\sum_{j}(JZ_{j}-hZ_{j}Z_{j+1})/\sqrt{L}\right)$ in (b,d), $\gamma=0.75$. In (c) the data with $L=6,8,10$ are almost identical. The inset shows $J_{L=10}/J_{L=8}$.
The above argument can be adapted to any weak measurements on a few sites, as the dissipative evolution $\mathcal{DL}$ will only preserve the identity operator generically. Numerically we observe that $J_{t}$ decays exponentially, see Fig. 2-(a). An example of a non-generic dynamics is spatial translation. It moves the operator away from the dephasing sites. In this case the dynamical entropy rate is positive, and related to the thermal entropy density [10]. Another exception is when $K_{m}$ is conserved by the dynamics. Then, though $J_{t}\not\to 0$ , the outcome Shannon entropy will not grow, and the dynamical entropy rate still vanishes.
To recapitulate, microscopic measurements are in general too intrusive to yield a positive entropy rate. Therefore we shall focus on the alternative of measuring a “mesoscopic observable” of the form
$Q=\frac{1}{\sqrt{V}}\sum_{x}Q_{x}\,,\,\left<Q_{x}\right>=0\,$, (10)
where $\left<A\right>:=\mathrm{Tr}[\rho A]$, $Q_{x}$ is a local operator centered around $x$, and $V$ is the volume (number of sites in a lattice model). We can always have $\left<Q_{x}\right>=0$ upon replacing $Q_{x}\to Q_{x}-\left<Q_{x}\right$>, so that $Q$ describes the fluctuation of an extensive quantity around its ensemble average (hence the term “mesoscopic”). Such coarse-grained quantities are known to behave classically, in the sense that their monitoring causes vanishing disturbance in the thermodynamic limit [63, 64]. Even in tiny systems, we observe numerically that the decay of $J_{t}$ slows down as the volume increases, see Fig. 2-(b). In the thermodynamic limit, we expect $J_{t}$ to reach a nonzero plateau from analytical calculation, see below.
Entropy rate formula. To facilitate analytical progress, we consider a general class of mesoscopic monitoring with continuous-valued outcome (replacing a sum over outcome by an integral):
$K_{m}=\frac{1}{(2\pi)^{\frac{1}{4}}}e^{-\left(Q\sqrt{\delta t}-m\right)^{2}/4}\,,\,m\in\mathbb{R}\,$, (11)
where $Q$ is as in (10). The $\sqrt{\delta t}$ factor allows a well-defined continuous monitoring $\delta t\to 0$ limit. The measurement strength can be tuned by rescaling $Q$ itself.
In the thermodynamic limit and away from any thermal criticality, the time-dependent fluctuation of $Q$ becomes Gaussian due to cluster decomposition (see Appendix). As a consequence the dynamical entropy rate can be computed solely in terms of the Keldysh and retarded Green functions in frequency domain [65]. We define them with the following convention (we assume $Q$ to be bosonic):
$G_{K}(\omega)=\frac{1}{2}\int\left<QQ(t)+Q(t)Q\right>e^{i\omega t}dt,$ (12a)
$G_{R}(\omega)=\frac{i}{2}\int\theta(t)\left<[Q(t),Q]\right>e^{i\omega t}dt.$ (12b)
At thermal equilibrium, they are related by the fluctuation-dissipation theorem (FDT), $\mathrm{Im}G_{R}(\omega)=\tanh(\beta\omega)G_{K}(\omega)/2$; $\mathrm{Re}G_{R}(\omega)$ is given by Kramers-Kronig. In terms of the eigenstate thermalization hypothesis [66, 67, 25], the Green functions encode the same data as the two-point structure factor (however the Gaussianity is not equivalent to the vanishing of higher free cumulants [68, 69]).
The Gaussianity enables a simple method to compute the dynamical entropy. We write a free boson model in which a displacement operator has the same $G_{K}$, and perform exact calculation therein (see Appendix). The result is the following:
$s_{\text{CNT}}:=\lim_{\begin{subarray}{c}t\to\infty\\ \delta t\to 0\end{subarray}}\lim_{V\to\infty}(S_{\text{CNT}}/t)$ (13)
$=$ $\frac{1}{2}\int\frac{d\omega}{2\pi}\ \left[\ln(1+\tilde{G})-\tilde{G}+\frac{\beta\omega}{\sinh\beta\omega}G_{K}\right]$
where $\tilde{G}(\omega):=G_{K}(\omega)+|G_{R}(\omega)|^{2}$. The three terms in the integral correspond to those in the definition (Planckian bound on quantum dynamical entropy), respectively, if the Shannon entropy is “renormalized”, that is, with the $Q=0$ value subtracted. As sanity checks, we observe the subadditivity of Shannon entropy from $\ln(1+\tilde{G})\leq\tilde{G}$, and the positivity of the entropy defect using $G_{K}{\beta\omega}/{\sinh\beta\omega}\leq G_{K}+|G_{R}|^{2}$.
Figure 3: Keldysh correlation function $G_{K}$ (as function of $\beta\omega$) that maximizes $\beta s_{\text{CNT}}$ (13). $G_{K}(-\omega)=G_{K}(\omega)$, obtained from numerically maximizing (13) subject to the fluctuation-dissipation theorem. Inset: comparison with $4/|\beta\omega|^{-1}$ (dashed line).
The entropy rate given by (13) satisfies a Planckian bound. To see this, we make a change of variable $y=\beta\omega$ in its right hand side, which then equals $\beta^{-1}$ times a $\beta$-independent integral, if we view $G_{K}$ and $G_{R}$ as functions of $y$. An explicit but untight upper bound can be obtained by maximizing the integral with respect to $G_{K}$ and $G_{R}$ ignoring FDT. This gives the optimum $|G_{R}|^{2}=0,G_{K}=y/(\sinh y-y)$, and
$\beta s_{\text{CNT}}\leq\int\frac{dy}{4\pi}\left[\ln\frac{\sinh(y)}{\sinh(y)-y}-\frac{y}{\sinh y}\right]\approx 0.57$ (14)
We may obtain a numerically tight bound by solving the FDT-constrained maximization: $\beta s_{\text{CNT}}\leq 0.375\dots$. The optimal $G_{K}$ (see Fig. 3) has a singularity $G_{K}\sim|\beta\omega|^{-\alpha}$ with $\alpha\approx 1$ at $\omega\to 0$.
The above results indicate that, at low temperature, the dominant contribution to a large dynamical entropy rate comes from the Planckian time scale ($\omega\sim 1/\beta$). Meanwhile high frequency contribution is suppressed due to the factor $\beta\omega/\sinh\beta\omega$ in (13). Intuitively, this is because correlation functions across the thermofield double of the form $\mathrm{Tr}[O\sqrt{\rho}Q(t)\sqrt{\rho}]$ are suppressed at $|\omega|\gtrsim\beta$, as they must decay as $e^{-\beta|\omega|/2}$ or faster [31, 40], and the information on $\bar{A}$ accessible from $A$ is encoded in such correlation functions. This effect is not artificial: the purification of the thermal initial state is essentially unique, and is necessary to encode information on initial condition (see [70, 71, 72, 73, 74] for parallel discussion on OTOCs).
We conjecture that the dynamical entropy rate obeys a universal Planckian bound $\beta s_{\text{CNT}}\leq c$, with possibly a different $c$ and the exception of thermal critical points, for any “single-operator” monitoring. To support this conjecture, recall first that microscopic monitoring generically yields vanishing entropy. An explicit Planckian bound is shown for the class of mesoscopic continuous monitoring (10)-(11); our calculation clearly also applies when a portion of the system is monitored. By interpolation, the bound should hold however the monitored region scales with system size. Since a large entropy rate is dominated by Planckian time scale fluctuations, changing details of the monitoring scheme [such as $\delta t$ or $K_{m}$ (11)] can at most modify the constant $c$ by a finite amount.
The rate of purification (4) under the same mesoscopic monitoring (10)-(11) can be computed with the same assumptions and method as (13) (see Appendix):
$\mathcal{J}:=\lim(J/t)=\frac{1}{2}\int\frac{d\omega}{2\pi}\frac{\beta\omega}{\sinh\beta\omega}\frac{G_{K}}{1+G_{K}+|G_{R}|^{2}}$ (15)
where the limit is the same as in (13). Interestingly, the purification rate (15) also satisfies a Planckian bound, thanks to a similar high-frequency suppression. Since both $G_{K}$ and $|G_{R}|^{2}$ are positive, we have $G_{K}/(1+G_{K}+|G_{R}|^{2})\leq 1$ implying
$\beta\mathcal{J}\leq\pi/8\approx 0.392\,.$ (16)
This bound is derived for a mesoscopic observable. For microscopic ones, we observe numerically that $J/t$ barely depends on system size, see Fig. 2-(c), indicating that it should vanish in the $V\to\infty,t\to\infty$ limit. Hence we conjecture that single-measurement purification also satisfies a general Planckian bound. By contrast, in Fig. 2-(d), we observe that a mesoscopic $J/t$ has a pronounced finite-size dependence similar to $J_{t}$, pointing to the direction of a nonzero asymptotic in the thermodynamic limit. Quantitative checks of the rate formulas will require efficient numerical methods [75, 76] and are left to future work.
Discussion. We close with some perspectives. Our results concern the “initial” time regime, where $t$ is large compared to the inverse coupling but small compared to any size-dependent time scales (such as the Thouless time). In this regime the system is effectively non-interacting for the coarse-grained monitoring (10), and the entropy rate is not directly related to microscopic features such as quantum integrability. It would be interesting to understand contributions from nonlinearity at later times and make connection with hydrodynamics [7, 77, 78, 79, 80, 81, 81, 82].
We focused on systems away from semiclassical or large $N$ limit. Our results are not useful for semiclassics, however they do apply to large $N$ models like SYK (see [83] for differences between semiclassics and large $N$), where a comparison with OTOCs would be interesting. More generally we would like to identify low-temperature states with large dynamical entropy. Finally a challenge is to lift the “single-operator” restriction in a sensible way [84], and propose a quantum many-body extension of the Pesin relation [17, 85, 86, 87].
Acknowledgements.
I thank Denis Bernard, Jorge Kurchan and Tomaž Prosen for sharing their insight on related topics and encouraging me to pursue this work.